A library take charge of rs. 5 for issue a book for one day and rs.1 thereafter. If ritu has taken 2 books for x days and y be the total amount of paid.write the linear variable in two variable for the situation. Plot the graph and find the amount to be paid for 5 days
1
Expert's answer
2019-12-27T09:01:12-0500

Given, y be the total amount of paid and x be the number of days.

Then, linear equation in two variables is:

"y=2[5 \times 1+1 \times (x-1)]"

"y=2[5 +(x-1)]"

"y=2[x+4]"

"y=2x+8"

For plotting the graph, taking two points.

Put "x=0"

"y=2 \times 0+8"

"y=8"

We get "(0,8)" .

Next, put "x=1"

"y=2 \times 1+8"

"y=10"

We get "(1,10)"

Plotting these points on the graph and joining them, we get a straight line as shown.



Next, for 5 days, put "x=5"

"y=2 \times5+8"

"y=18"

So, total amount paid for 5 days is Rs. 18.